An isosceles triangles's base is 15 cm and the other two sides are 21cm. Find its area
Answer by Boreal(15235)

The isosceles triangle can have the base split into two halves, each 7.5 cm long.
Now the altitude is one of the legs of a right triangle with the other leg 7.5 and hypotenuse 21
56.25 (1/2 of base squared) + x^2 = 441 (21^2) Pythagorean Theorem.
Subtract 56.25 from each side.
x^2= 384.75
x=19.61 cm positive root only.
That is the altitude.
The area is (1/2) bh
base is 15 cm, altitude 19.61 cm
0.5 * 15 *19.61=147.075 cm^2